sin10.sin30.sin50.sin70.sin90=?
what should be the answer for this difficult question please tellme

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As sin 90 = 1
We have sinx*sin(60 – x)*sin(60 + x) = 1/4* sin3x
sin10*sin30*sin50*sin70
=1/2*sin10*sin(60 – 10)*sin(60 + 10)
=1/2*1/4*sin(3*10)
=1/2*1/4*1/2
=1/16